896 mL of an ideal gas at 760 torr is compressed to 561 mL at constant temperature. What is the new pressure?
Answer #1

P1V1 = P2V2

where, P1 = 760 torr = 1 atm

V1 = 896 mL = 0.896 L

P2 = ?

V2 = 561 mL = 0.561 L

1 * 0.896 = P2 * 0.561

0.896 = P2 * 0.561

P2 = 0.896 / 0.561 = 1.60 atm

Therefore, the new pressure = 1.60 atm